# MIT-CTP/5802

### v2

@@ 1 Introduction
Our goal is to provide a way to systematically exploit Lorentz equivariance in different networks tasks defined on relativistic phase space. A Lorentz-equivariant transformer can provide the appropriate internal or latent representation of phase space data to regression, classification, and generative networks. L-GATr is based on the Geometric Algebra Transformer (GATr) [33, 34], designed for Euclidean translations, rotations, and reflections. We generalize it to L-GATr encoding Lorentz-equivariance into new network layers, including a maximally expressive linear map, attention, and layer normalization. This architecture was originally developed for an ML audience and applied to amplitude regression, top tagging, and event generation in Ref. [35]. In this physics-targeted study, we extend the amplitude regression part, improve the classification through pre-training and multi-class tagging, and deliver a competitive generative network. We also provide a comprehensive benchmarking with the state of the art for these three distinct LHC applications.
After reviewing the construction of L-GATr in Sec. 2, we demonstrate the versatility of L-GATr through three LHC case studies. In Sec. 3 we show how it allows us to efficiently learn precision surrogates for scattering amplitudes up to a $Z+5$ gluon final state. Next, we show in Sec. 4 how we can improve transformer-based jet taggers with an equivariant setup. We show how L-GATr benefits from pre-training for the ultimate performance and can be generalized to multi-class tagging. Finally, in Sec. 5 we employ L-GATr inside a diffusion generator and show how it generates LHC events for instance for final states up to $t\bar{t}+4$ jets better than all benchmarks. We provide a brief summary and outlook in Sec. 6. Throughout the text and for all figures we indicate possible overlap with Ref. [35].
@@ 2 Lorentz-Equivariant Geometric Algebra Transformer
In this section, we discuss the principles behind L-GATr and its most important features. L-GATr uses the geometric algebra, a mathematical framework that represents certain geometric objects and operations in a unified language. Using the language of geometric algebra, it is straightforward to build equivariant layers while retaining most of the structure from typical neural networks. In cases where the Lorentz symmetry is not completely realized, we show how reference inputs can be used to make L-GATr equivariant with respect to specific subgroups of the Lorentz group. Finally, we study how L-GATr scales with the phase space dimensionality, as compared to other network architectures.
@@ 2.1 Spacetime Geometric Algebra
where the anti-commutator $\{x,y\}/2$ represents the usual inner product, and $[x,y]/2$ constitutes a new outer product. This second term defines the bivector, which can be understood geometrically as an area element of the plane spanned by $x$ and $y$. The geometric product $xy$ is an element of the algebra made up by the sum of a scalar and a bivector. Neither object is part of the original vector space, so the geometric product extends the original vector space. This extension can be carried out systematically by applying the geometric product repeatedly to the basis elements of the vector space. To develop L-GATr, we focus on the etric alg $\mathbb{G}_{1,3}$, built from the vector space $\mathbb{R}^{4}$ with the metric $g=\text{diag}(1,-1,-1,-1)$. We choose as a basis for the vector space a set of four real vectors $\gamma^{\mu}$, which satisfy the anti-commutation relation
@@
This inner product establishes the basis elements as a set of orthogonal vectors and fixes their normalization. This happens to also be the defining property of the gamma matrices, the basis elements of the Dirac algebra used to describe spinor interactions. Both algebras are closely related, the only difference being that the spacetime algebra is defined over $\mathbb{R}^{4}$, whereas the Dirac algebra is defined over $\mathbb{C}^{4}$. This prescription fully recovers all the algebra properties presented in Ref. [35].
We now construct new elements of the algebra using the geometric product defined in Eq. (1). All higher-order elements can be characterized as antisymmetric products of $\gamma^{\mu}$. We organize them in grades, defined by the number of $\gamma^{\mu}$ needed to express them. For instance, the antisymmetric tensor $\sigma^{\mu\nu}$ is generated from the geometric product of two $\gamma^{\mu}$ and consequently has grade two,
@@
Following Eq. (1), $\sigma^{\mu\nu}$ is a bivector, which can be interpreted as the plane opened by $\mu$ and $\nu$ in Minkowski space. We see that the symmetric term in the geometric product reduces the grade, while the antisymmetric term increases it. The whole product $\gamma^{\mu}\gamma^{\nu}$ is a sum of grade zero (scalar) and grade two. A generic element of the algebra that mixes grade information is called a multivector.
Moving on, the geometric product of three vectors $\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}$ contains the antisymmetric tensor $\epsilon_{\mu\nu\rho\sigma}\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}$ as a trivector or axial vector. The product of all four $\gamma^{\mu}$ leads us to the pseudoscalar
$\gamma^{5}=\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}\equiv\frac{1}{4!}\epsilon_{\mu\nu\rho\sigma}\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}\gamma^{\sigma}.$ (4)
Pseudoscalars act as parity reversal operations on any object and can be used to write axial vectors as $\gamma^{\mu}\gamma^{5}$. The missing factor $i$ compared to the usual definition of $\gamma^{5}$ indicates the slight difference between the complex Dirac algebra and our real spacetime algebra.
@@
In this representation, we only include the nonzero and independent entries in the antisymmetric bivector. Multivectors can be used to represent both spacetime objects and Lorentz transformations. For instance, particles are characterized by their type (i.e. particle identification, or PID) and their 4-momentum $p^{\mu}$,
@@
Using this convenient representation, the spacetime algebra naturally structures relevant objects like parity-violating transition amplitudes. The matrix element $\mathcal{M}$ is a function of 4-momenta and can be decomposed into parity-even and parity-odd terms before it gets squared,
@@
The first two terms represent a scalar function of the 4-momenta, while the last is a pseudoscalar function. To calculate this amplitude in the spacetime algebra, we first embed each 4-momentum into a multivector $x^{i}=p_{\mu}^{i}\gamma^{\mu}$. Using these multivectors as inputs, the squared amplitude can be obtained through a sequence of algebra operations. The result of this calculation will also be a multivector, namely
@@
The geometric algebra explicitly separates the scalar and pseudoscalar components of the squared amplitude, highlighting their respective geometric significance.
The geometric algebra also allows us to perform operations on spacetime objects. Lorentz transformations act as
@@
where $v$ is a multivector representing an element of the Lorentz group acting on the algebra element $x$, and $v^{-1}$ representings the corresponding inverse. The representation $v$ of a Lorentz transformation is built by a simple rule: a multivector encoding an object that is invariant under a Lorentz transformation will also represent the transformation itself. This gives a dual interpretation to spacetime algebra elements as, both, geometric objects and Lorentz transformations.
@@
$v=e^{\omega\sigma^{03}}=1\cosh\omega+\sigma^{03}\sinh\omega\;.$ (10)
@@ 2.2 Constructing a Lorentz-Equivariant Architecture
Based on the multivector representation, we now construct the corresponding transformer network L-GATr. It is equivariant under Lorentz group transformations $\Lambda$
[…]
We take advantage of the fact that multivector grades form sub-representations of the Lorentz group, i.e. all multivector components of the same grade transform equally under all network operations, whereas different grades transform differently [33, 35].
[…]
@@
[…]
@@
  • For the linear layers, we use the fact that equivariant operations on multivectors process components within the same grade equally. We use the projection $\langle\cdot\rangle_{k}$ to extract the $k$-grade and apply different learnable coefficients for each grade. As a result, the most general linear combination of independently-transforming multivector components is
    $\text{Linear}(x)=\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}\quad\left(+\sum_{k=0}^{4}\,w_{k}\gamma^{5}\langle x\rangle_{k}\right)\;,$ (15)
@@
  • We extend scaled dot-product attention such that it can be applied to multivectors
@@
[…]
@@
    where $\epsilon$ is a normalization constant and $n_{c}$ is the number of multivector channels.
  • Activation functions applied directly on the multivectors break the equivariance. We employ scalar-gated activation functions [33], where the nonlinearity only acts on the scalar component of the multivector $\langle x\rangle_{0}$. Specifically, we use the scalar-gated GELU [40] activation function
@@
$\text{Linear}(x)$ $vx+w$ $\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}$
@@ 2.3 Breaking Lorentz Symmetry
In many LHC contexts, Lorentz symmetry is only partially preserved. L-GATr can apply partial symmetry breaking in a tunable manner by including reference multivectors as additional inputs. Any network operation that involves such a reference vector will violate equivariance, breaking the symmetry group to a subgroup where the reference direction is fixed. This defines a partial symmetry breaking without altering the structure of the network. The network always has the option to tune out the reference vectors when they are not needed. Reference vectors are appended for each L-GATr input $x$ in the same way, either as extra tokens, or as extra channels within each token.
For instance, the LHC beam direction breaks the Lorentz group to the subgroup of rotations around and boosts along the beam axis [41, 42, 14, 10]. The natural reference vector is this beam direction, which can be either implemented as two vectors $x^{V}_{\pm}=(0,0,0,\pm 1)$, or one bivector representing the $x-y$ plane, $x^{B}_{12}=1$. We find similar performance for both choices. Generally, we can break the Lorentz group to the subgroup of rotations in three-dimensional space $SO(3)$ using the reference multivector $x^{V}=(1,0,0,0)$.
@@ 2.4 Scaling with the Number of Particles
Figure 1: Scaling behavior of L-GATr, a standard transformer, and the equivariant graph network CGENN. The left panel was already discussed in Ref. [35].
Fully connected graph neural networks bear a very close resemblance to transformers [43]; both process data as sets of tokens, both respect full or partial permutation symmetry, and both can be turned equivariant [10, 12, 14, 35]. Resource efficiency is where the two architectures differ most. To quantify it, we measure the scaling of speed and memory consumption of a standard transformer, L-GATr, and the Clifford Group Equivariant Neural Network (CGENN) [14], which is a graph network built on geometric algebra representations. We expect the CGENN to represent the strengths and limitations of equivariant graph networks. We execute network forward passes with synthetic data on an H100 GPU. Each measurement is done with a single event consisting of a varying number of particles. To ensure fairness, all networks consist of a single network block, for the transformers the attention receives inputs with 72 channels from 4 attention heads, and the CGENN is set up to be as close as possible to L-GATr. Namely, both contain 8 multivector channels and 16 scalar channels.
In the left panel of Fig. 1, we see that the evaluation time of all networks is independent of the number of tokens in the few-token regime, but it eventually scales quadratically. This transition happens when attention or message passing, rather than other parallelizable operations, becomes the limiting factor. L-GATr scales like a standard transformer in the many-token regime because they both use the same attention module. For few tokens, L-GATr is slower because of the more expensive linear layers. The CGENN is slower than L-GATr for few tokens, and the quadratic scaling due to the expensive message passing operation takes off sooner.
As can be seen in the right panel of Fig. 1, for many particles L-GATr and the standard transformer display the same linear scaling in memory usage with the number of tokens since they use the same attention module. In contrast, the CGENN scales quadratically in this regime and runs out of memory already for 1000 particles. We attribute this to the different degree of optimization in the architectures. For L-GATr we use FlashAttention [44], heavily optimized for speed and memory efficiency. Graph neural networks are often optimized for sparsely connected graphs, so the efficiency of the standard implementation degrades for fully connected graphs.

### v3

@@ 1 Introduction
Our goal is to provide a way to systematically exploit Lorentz equivariance in different network tasks defined on relativistic phase space. A Lorentz-equivariant transformer can provide the appropriate internal or latent representation of phase space data to regression, classification, and generative networks. L-GATr is based on the Geometric Algebra Transformer (GATr) [33, 34], designed for Euclidean translations, rotations, and reflections. We generalize it to L-GATr encoding exact Lorentz-equivariance into new network layers, including a maximally expressive linear map, attention, and layer normalization. This architecture was originally developed for an ML audience and applied to amplitude regression, top tagging, and event generation in Ref. [35]. In this physics-targeted study, we extend the amplitude regression analysis, improve the classification through pre-training and multi-class tagging, and deliver a competitive generative network for Monte Carlo event generation. We also provide a comprehensive benchmarking with the state of the art for these three distinct LHC applications.
After reviewing the construction of L-GATr in Sec. 2, we demonstrate the versatility of L-GATr through three LHC case studies. In Sec. 3 we show how it allows us to efficiently learn precision surrogates for scattering amplitudes up to a $Z+5$ gluon final state. Next, we show in Sec. 4 how we can improve transformer-based jet taggers with an equivariant setup. We show how L-GATr benefits from pre-training for the ultimate performance and can be generalized to multi-class tagging. Finally, in Sec. 5 we employ L-GATr inside a diffusion generator and show how it generates LHC events for final states up to $t\bar{t}+4$ jets better than all benchmarks. We provide a brief summary and outlook in Sec. 6. Throughout the text and for all figures we indicate possible overlap with Ref. [35].
@@ 2 Lorentz-Equivariant Geometric Algebra Transformer
In this section, we discuss the principles behind L-GATr and its most important features. L-GATr uses the geometric algebra, a mathematical framework that represents certain geometric objects and operations in a unified language. Using the language of geometric algebra, it is straightforward to build exactly equivariant layers while retaining most of the structure from typical neural networks. In cases where the Lorentz symmetry is not completely realized, we show how reference vector inputs can be used to make L-GATr equivariant with respect to specific subgroups of the Lorentz group. Finally, we study how L-GATr scales with the phase space dimensionality as compared to other network architectures.
@@ 2.1 Spacetime Geometric Algebra
where the anti-commutator $\{x,y\}/2$ represents the usual inner product, and $[x,y]/2$ constitutes a new outer product. This second term defines the bivector, which can be understood geometrically as an area element of the plane spanned by $x$ and $y$. Neither of the terms in the geometric product is an element of the original vector space, so the geometric product extends it into a larger domain with higher-order geometric objects.
To develop L-GATr, we focus on the spacetime algebra $\mathbb{G}_{1,3}$, built from the vector space $\mathbb{R}^{4}$ with the metric $g=\text{diag}(1,-1,-1,-1)$. We choose as a basis for the vector space a set of four real vectors $\gamma^{\mu}$, which satisfy the anti-commutation relation
@@
This inner product establishes the basis elements as a set of orthogonal vectors and fixes their normalization. This prescription fully recovers all the algebra properties presented in Ref. [35]. Futhermore, Eq. (2) is also the defining property of the gamma matrices, the basis elements of the Dirac algebra used to describe spinor interactions. Both algebras are closely related, the only difference being that the spacetime algebra is defined over $\mathbb{R}^{4}$, whereas the Dirac algebra is defined over $\mathbb{C}^{4}$.
We now construct the full range of algebra elements by using the geometric product defined in Eq. (1). All higher-order elements can be characterized as antisymmetric products of $\gamma^{\mu}$. We organize them in grades, defined by the number of $\gamma^{\mu}$ needed to express them. The first higher-order object is the antisymmetric tensor $\sigma^{\mu\nu}$, which is generated from the geometric product of two $\gamma^{\mu}$ and consequently has grade two,
@@
Following Eq. (1), $\sigma^{\mu\nu}$ is a bivector, which can be interpreted as the plane defined by $\mu$ and $\nu$ in Minkowski space. We see that the symmetric term in the geometric product reduces the grade, while the antisymmetric term increases it.
The next higher-order object emerges from the geometric product of three vectors $\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}$, which generates the trivector or axial vector in the form of the antisymmetric tensor $\epsilon_{\mu\nu\rho\sigma}\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}$. Finally, the last higher-order object in the algebra is the pseudoscalar
$\gamma^{5}=\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}\equiv\frac{1}{4!}\epsilon_{\mu\nu\rho\sigma}\gamma^{\mu}\gamma^{\nu}\gamma^{\rho}\gamma^{\sigma},$ (4)
which can be obtained through the product of all four $\gamma^{\mu}$. Pseudoscalars act as parity reversal operations on any object and can be used to write axial vectors as $\gamma^{\mu}\gamma^{5}$. As a side note, the missing factor $i$ compared to the usual definition of $\gamma^{5}$ indicates the slight difference between the complex Dirac algebra and the real spacetime algebra.
@@
In this representation, we only include the nonzero and independent entries in the bivector. We call this object a multivector, which is the name for a general algebra element. Multivectors can be used to represent both spacetime objects and Lorentz transformations. We want to apply this representation to particles, which can be characterized by their type (i.e. particle identification, or PID) and their 4-momentum $p^{\mu}$,
@@
Using this representation, the spacetime algebra naturally structures relevant objects like parity-violating transition amplitudes. The matrix element $\mathcal{M}$ is a function of 4-momenta and can be decomposed into parity-even and parity-odd terms before it gets squared,
@@
The first two terms represent a scalar function of the 4-momenta, while the last is a pseudoscalar function. To calculate this amplitude in the spacetime algebra, we first embed each 4-momentum into a multivector $x^{i}=p_{\mu}^{i}\gamma^{\mu}$. Working from this representation, the squared amplitude can be obtained through a sequence of algebra operations as
@@
This result is also a multivector, which explicitly separates the scalar and pseudoscalar components of the squared amplitude.
The geometric algebra also allows us to apply Lorentz transformations on spacetime objects as
@@
where $v$ is a multivector representing an element of the Lorentz group acting on the algebra element $x$, and $v^{-1}$ is its corresponding inverse. The representation $v$ of a Lorentz transformation is built by a simple rule: a multivector encoding an object that is invariant under a Lorentz transformation will also represent the transformation itself. This gives a dual interpretation to spacetime algebra elements as both geometric objects and Lorentz transformations.
@@
$v=e^{\omega\sigma^{03}/2}=1\cosh\frac{\omega}{2}+\sigma^{03}\sinh\frac{\omega}{2}\;.$ (10)
@@ 2.2 Constructing a Lorentz-Equivariant Architecture
Based on the multivector representation, we now construct the transformer network L-GATr. It is exactly equivariant under Lorentz group transformations $\Lambda$
$\text{L-GATr}\Big{(}\Lambda(x)\Big{)}=\Lambda\Big{(}\text{L-GATr}(x)\Big{)}\;.$ (12)
To ensure that this property is preserved, we take advantage of the fact that multivector grades form sub-representations of the Lorentz group, i.e., all multivector components of the same grade transform equally under all network operations, whereas different grades transform differently [33, 35].
The L-GATr architecture uses variations of the standard transformer operations Linear, Attention, LayerNorm, and Activation, adapted to process multivectors [38, 39]. As usual for transformers, the input $x$ and output $\text{L-GATr}(x)$ are unordered sets of $n_{t}$ ordered lists of $n_{c}$ multivector channels
@@
$\text{L-GATr}(x)$ $=\text{Linear}\circ\text{Block}\circ\text{Block}\circ\cdots\circ\text{Block}\circ\text{Linear}(x)\;.$ (14)
@@
  • For the linear layers, we use the fact that equivariant operations on multivectors process components within the same grade equally. We use the projection $\langle\cdot\rangle_{k}$ to extract the $k$-th grade and apply different learnable coefficients for each grade. As a result, the most general linear combination of independently-transforming multivector components is
    $\text{Linear}(x)=\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}\quad+\sum_{k=0}^{4}\,w_{k}\gamma^{5}\langle x\rangle_{k}\;,$ (15)
@@
  • We extend the scaled dot-product attention such that it can be applied to multivectors
@@
    where $n_{c}$ is the number of multivector channels and $\langle\cdot,\cdot\rangle$ is the $\mathbb{G}_{1,3}$ inner product. This inner product can be pre-computed as a list of signs and a Euclidean inner product, allowing us to use standard transformer implementations.
@@
    where $\epsilon=10^{-2}$ is a normalization constant and $n_{c}$ is the number of multivector channels. We find that the specific value of $\epsilon$ has a negligible impact on the performance of the network.
  • Activation functions applied directly on the multivectors break the equivariance. We employ scalar-gated activation functions [33], where the whole multivector is scaled by a nonlinearity acting only on the scalar component of the multivector $\langle x\rangle_{0}$. Specifically, we use the scalar-gated GELU [40] activation function
@@
$\text{Linear}(x)$ $vx+w$ $\sum_{k=0}^{4}v_{k}\langle x\rangle_{k}+\sum_{k=0}^{4}w_{k}\gamma^{5}\langle x\rangle_{k}$
@@ 2.3 Breaking Lorentz Symmetry
In many LHC contexts, Lorentz symmetry is only partially preserved. If this partial symmetry breaking is not accounted for when applying a Lorentz-equivariant architecture, performance can degrade significantly. This is because these architectures inherently treat inputs related by global Lorentz transformations as equivalent. However, in the presence of symmetry breaking, these inputs may carry different physical information. A fully Lorentz-equivariant architecture is, by construction, blind to these differences, which can limit its effectiveness.
L-GATr can apply partial symmetry breaking in a tunable manner by including reference multivectors as additional inputs. Any network operation that involves such a reference vector will violate equivariance, breaking the symmetry group to a subgroup where the reference direction is fixed. This defines a partial symmetry breaking without altering the structure of the network. The network has the option to tune out the reference vectors in certain phase space regions or even globally if they are not required. As we discuss below, this strategy produces better results than a network where the symmetry is completely broken. Reference vectors are appended for each L-GATr input $x$ in the same way, either as extra tokens, or as extra channels within each token. We emphasize that the reference multivectors should be regarded as part of the network architecture rather than as input data components. For example, when applying data augmentation, one should augment only the input particle data, and then append the reference multivectors directly before processing the particles with the Lorentz-equivariant architecture.
There are two symmetry breaking sources that should be taken into account when processing reconstructed particles at the LHC. The first one is the LHC beam direction, which breaks the Lorentz group to the subgroup of rotations around and boosts along the beam axis [41, 42, 14, 10]. The natural reference vector is the beam direction itself, which can be either implemented as two vectors $x^{V}_{\pm}=(0,0,0,\pm 1)$, or one bivector representing the xyx-y plane, $x^{B}_{12}=1$. We find similar performance for both choices. The second source of symmetry breaking is the detector setup, which can compromise the symmetry of the observables with respect to relativistic boosts. Generally, we can break the Lorentz group to the subgroup of rotations in three-dimensional space $SO(3)$ using the reference multivector $x^{V}=(1,0,0,0)$. A pair of multivectors $x^{V}_{\pm}=(1,0,0,\pm 1)$, which represent the beam with a non-zero time component, also break full Lorentz equivariance down to $\mathrm{S}O(2)$ equivariance under rotations around the beam axis. This reduction occurs because these multivectors are linear combinations of vectors pointing in the time and beam directions.
@@ 2.4 Scaling with the Number of Particles
Figure 1: Scaling behavior of L-GATr, a standard transformer, and the equivariant graph network CGENN. The left panel was already discussed in Ref. [35]. Error bands are based on the mean and standard deviation from 3 separate evaluations and originate from the GPU randomness and the time measuring software.
Fully connected graph neural networks bear a very close resemblance to transformers [43]; both process data as sets of tokens, both respect full or partial permutation symmetry, and both can be turned equivariant [10, 12, 14, 35]. Resource efficiency is where the two architectures differ most. To quantify it, we measure the scaling of speed and memory consumption of a standard transformer, L-GATr, and the Clifford Group Equivariant Neural Network (CGENN) [14], which is a graph network built on geometric algebra representations. We expect the CGENN to represent the strengths and limitations of equivariant graph networks. We perform the measurements on an H100 GPU by evaluating the network on datasets containing a single event and a varying number of particles. For this test we use Gaussian noise to build the datasets and do not train the network, since we are only concerned with the resource management of the architectures. All networks consist of a single network block. For both transformer-based architectures, the attention receives inputs with 72 channels from 4 attention heads. In the case of L-GATr, we achieve this by building the architecture with 8 multivector channels and 16 scalar channels. This configuration is chosen because the attention operation represents the main computational bottleneck for both networks, and we want to ensure a fair comparison of its contribution during the measurements. As for CGENN, it is set up to be as close as possible to L-GATr, featuring the same hidden channel structure. In total, L-GATr consists of $2.3\times 10^{4}$ parameters, the transformer consists of $6.7\times 10^{5}$ parameters, and CGENN consists of $2.5\times 10^{4}$ parameters.
In the left panel of Fig. 1, we see that the evaluation time of all networks is independent of the number of tokens in the few-token regime, but it eventually scales quadratically. This transition happens when attention or message passing, rather than other parallelizable operations, becomes the limiting factor. L-GATr scales like a standard transformer in the many-token regime because they both use the same attention module. For few tokens, L-GATr is slower because of the more expensive linear layers. CGENN is slower than L-GATr for few tokens, and the quadratic scaling due to the expensive message passing operation takes off sooner.
As can be seen in the right panel of Fig. 1, for many particles L-GATr and the standard transformer display the same linear scaling in memory usage with the number of tokens since they use the same attention module. In contrast, CGENN scales quadratically in this regime and runs out of memory already for 1000 particles. We attribute this to the different degree of optimization in the architectures. For L-GATr we use FlashAttention [44], heavily optimized for speed and memory efficiency. Graph neural networks are often optimized for sparsely connected graphs, so the efficiency of the standard implementation degrades for fully connected graphs.